normal body temperature is98.6 degress F . during an illness it rose 3.2 degressF . what was the new temperature?
Found 2 solutions by oberobic, jim_thompson5910:
Answer by oberobic(2304)

98.6 + 3.2 = 101.8
.

Answer by jim_thompson5910(35256)

New temperature = 98.6 + 3.2 = 101.8 degrees F